### v1

@@ Abstract
11footnotetext: Equal contributions22footnotetext: Formerly affiliated with misraj.ai at the time of this research; omar.hadid@gmail.com 33footnotetext: Kuwain: \<كُوَيْن>: is a diminutive form of the Arabic word (Kawn), which means "universe". So, "Kuwain"means "tiny universe"or "little cosmos".Kuwain is one of a series of Arabic-English multilingual Large Language Models (LLMs), embodying the concept of a vast amount of knowledge condensed into a compact, accessible form, much like a miniature universe of information.
@@ 1 Introduction
Рис. 1: Model performance vs size visualization: Points represent models on the Arabic NLP benchmark, where point size indicates efficiency (average score divided by model size). Larger points show better performance-to-size ratio.
@@ 4 Method
Рис. 2: (a) Vocabulary expansion by adding new Arabic tokens to the tokenizer. (b) Extension of model layers, where newly added layers were trainable while the original layers remained frozen.
@@ 4.1 Extending Model’s Layers
Рис. 3: models decoder architecture.
@@ 5.1 Number and Placement of New Layers
Рис. 4: Training loss charts for Kuwain under different configurations: (a) Freezing the last layer leads to instability, (b) Enabling the last layer and avoiding consecutive layers yields stable training, (c) Stacking new layers consecutively results in degraded performance.
@@ 5.2 Tokenizer Comparison
Таблица 1: Comparison of tokenization metrics across different Arabic language models. The vocab size used presents the Arabic vocab for each model, where the full vocab has a different number.
@@ 5.3 Data Proportion
Таблица 2: Performance comparison on English benchmarks between TinyLlama and two variants of Kuwain: trained with 20% English data (Kuwain) and with less than 20% English data (Kuwain-lt-$\phi$). Results show how English performance varies with data proportion.
@@ 5.4 Conventional Training (Without Layer Extension)
Таблица 3: English evaluation of TinyLlama vs. Kuwain-Naive (standard pre-training).
@@
Таблица 4: Arabic evaluation of Kuwain-Naive vs. our model Kuwain.
@@ 6 Evaluation
Таблица 5: Arabic Benchmark Evaluation of Kuwain and TinyLlama
@@
Таблица 6: Arabic Leader board evaluation
@@ 7 Conclusion
Список литературы
  • [1] Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. Llama 2: Open foundation and fine-tuned chat models. arXiv preprint arXiv:2307.09288, 2023.
  • [2] Xiang Zhang, Senyu Li, Bradley Hauer, Ning Shi, and Grzegorz Kondrak. Don’t trust chatgpt when your question is not in english: A study of multilingual abilities and types of llms. arXiv preprint arXiv:2305.16339, 2023.
  • [3] Viet Dac Lai, Nghia Trung Ngo, Amir Pouran Ben Veyseh, Hieu Man, Franck Dernoncourt, Trung Bui, and Thien Huu Nguyen. Chatgpt beyond english: Towards a comprehensive evaluation of large language models in multilingual learning. arXiv preprint arXiv:2304.05613, 2023.
  • [4] Albert Q Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra Singh Chaplot, Diego de las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, et al. Mistral 7b. arXiv preprint arXiv:2310.06825, 2023.
  • [5] Marah Abdin, Sam Ade Jacobs, Ammar Ahmad Awan, Jyoti Aneja, Ahmed Awadallah, Hany Awadalla, Nguyen Bach, Amit Bahree, Arash Bakhtiari, Harkirat Behl, et al. Phi-3 technical report: A highly capable language model locally on your phone. arXiv preprint arXiv:2404.14219, 2024.
  • [6] Gemma Team, Thomas Mesnard, Cassidy Hardin, Robert Dadashi, Surya Bhupatiraju, Shreya Pathak, Laurent Sifre, Morgane Rivière, Mihir Sanjay Kale, Juliette Love, et al. Gemma: Open models based on gemini research and technology. arXiv preprint arXiv:2403.08295, 2024.
  • [7] Aaron Grattafiori, Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Alex Vaughan, et al. The llama 3 herd of models. arXiv preprint arXiv:2407.21783, 2024.
  • [8] Peiyuan Zhang, Guangtao Zeng, Tianduo Wang, and Wei Lu. Tinyllama: An open-source small language model. arXiv preprint arXiv:2401.02385, 2024.
  • [9] M Ali, M Fromm, K Thellmann, R Rutmann, M Lübbering, J Leveling, K Klug, J Ebert, N Doll, JS Buschhoff, et al. Tokenizer choice for llm training: Negligible or crucial? preprint (2023). arXiv preprint arXiv:2310.08754, 2023.
  • [10] Mohamed Motasim Hamed, Muhammad Hreden, Khalil Hennara, Zeina Aldallal, Sara Chrouf, and Safwan AlModhayan. Lahjawi: Arabic cross-dialect translator. In Proceedings of the 4th Workshop on Arabic Corpus Linguistics (WACL-4), pages 12–24, 2025.
  • [11] Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, et al. Palm: Scaling language modeling with pathways. Journal of Machine Learning Research, 24(240):1–113, 2023.
  • [12] Teven Le Scao, Angela Fan, Christopher Akiki, Ellie Pavlick, Suzana Ilić, Daniel Hesslow, Roman Castagné, Alexandra Sasha Luccioni, François Yvon, Matthias Gallé, et al. Bloom: A 176b-parameter open-access multilingual language model. https://inria.hal.science/hal-03850124, 2023.
  • [13] L Xue. mt5: A massively multilingual pre-trained text-to-text transformer. arXiv preprint arXiv:2010.11934, 2020.
  • [14] Terra Blevins and Luke Zettlemoyer. Language contamination helps explain the cross-lingual capabilities of english pretrained models. arXiv preprint arXiv:2204.08110, 2022.
  • [15] Jared Kaplan, Sam McCandlish, Tom Henighan, Tom B Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei. Scaling laws for neural language models. arXiv preprint arXiv:2001.08361, 2020.
  • [16] Tom B Brown. Language models are few-shot learners. arXiv preprint arXiv:2005.14165, 2020.
  • [17] Uri Shaham, Jonathan Herzig, Roee Aharoni, Idan Szpektor, Reut Tsarfaty, and Matan Eyal. Multilingual instruction tuning with just a pinch of multilinguality. arXiv preprint arXiv:2401.01854, 2024.
  • [18] Kshitij Gupta, Benjamin Thérien, Adam Ibrahim, Mats L Richter, Quentin Anthony, Eugene Belilovsky, Irina Rish, and Timothée Lesort. Continual pre-training of large language models: How to (re) warm your model? arXiv preprint arXiv:2308.04014, 2023.
  • [19] Yiming Cui, Ziqing Yang, and Xin Yao. Efficient and effective text encoding for chinese llama and alpaca. arXiv preprint arXiv:2304.08177, 2023.
  • [20] Björn Plüster. Leolm: Igniting german-language llm research, 2023.
  • [21] Naoaki Okazaki, Kakeru Hattori, Hirai Shota, Hiroki Iida, Masanari Ohi, Kazuki Fujii, Taishi Nakamura, Mengsay Loem, Rio Yokota, and Sakae Mizuki. Building a large japanese web corpus for large language models. arXiv preprint arXiv:2404.17733, 2024.
  • [22] Leon Voukoutis, Dimitris Roussis, Georgios Paraskevopoulos, Sokratis Sofianopoulos, Prokopis Prokopidis, Vassilis Papavasileiou, Athanasios Katsamanis, Stelios Piperidis, and Vassilis Katsouros. Meltemi: The first open large language model for greek. arXiv preprint arXiv:2407.20743, 2024.
  • [23] Terra Blevins, Tomasz Limisiewicz, Suchin Gururangan, Margaret Li, Hila Gonen, Noah A Smith, and Luke Zettlemoyer. Breaking the curse of multilinguality with cross-lingual expert language models. arXiv preprint arXiv:2401.10440, 2024.
  • [24] Hao Zhou, Zhijun Wang, Shujian Huang, Xin Huang, Xue Han, Junlan Feng, Chao Deng, Weihua Luo, and Jiajun Chen. Moe-lpr: Multilingual extension of large language models through mixture-of-experts with language priors routing. arXiv preprint arXiv:2408.11396, 2024.
  • [25] Zheng-Xin Yong, Hailey Schoelkopf, Niklas Muennighoff, Alham Fikri Aji, David Ifeoluwa Adelani, Khalid Almubarak, M Saiful Bari, Lintang Sutawika, Jungo Kasai, Ahmed Baruwa, et al. Bloom+ 1: Adding language support to bloom for zero-shot prompting. arXiv preprint arXiv:2212.09535, 2022.
  • [26] Neha Sengupta, Sunil Kumar Sahu, Bokang Jia, Satheesh Katipomu, Haonan Li, Fajri Koto, Osama Mohammed Afzal, Samta Kamboj, Onkar Pandit, Rahul Pal, et al. Jais and jais-chat: Arabic-centric foundation and instruction-tuned open generative large language models. arXiv preprint arXiv:2308.16149, 2023.
  • [27] Huang Huang, Fei Yu, Jianqing Zhu, Xuening Sun, Hao Cheng, Dingjie Song, Zhihong Chen, Abdulmohsen Alharthi, Bang An, Ziche Liu, et al. Acegpt, localizing large language models in arabic. arXiv preprint arXiv:2309.12053, 2023.
  • [28] Anis Koubaa, Adel Ammar, Lahouari Ghouti, Omar Najar, and Serry Sibaee. Arabiangpt: Native arabic gpt-based large language. arXiv preprint arXiv:2402.15313, 2024.
  • [29] M Saiful Bari, Yazeed Alnumay, Norah A Alzahrani, Nouf M Alotaibi, Hisham A Alyahya, Sultan AlRashed, Faisal A Mirza, Shaykhah Z Alsubaie, Hassan A Alahmed, Ghadah Alabduljabbar, et al. Allam: Large language models for arabic and english. arXiv e-prints, pages arXiv–2407, 2024.
  • [30] Thuat Nguyen, Chien Van Nguyen, Viet Dac Lai, Hieu Man, Nghia Trung Ngo, Franck Dernoncourt, Ryan A Rossi, and Thien Huu Nguyen. Culturax: A cleaned, enormous, and multilingual dataset for large language models in 167 languages. arXiv preprint arXiv:2309.09400, 2023.
  • [31] Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J. Liu. Exploring the limits of transfer learning with a unified text-to-text transformer. arXiv e-prints, 2019.
  • [32] Chengyue Wu, Yukang Gan, Yixiao Ge, Zeyu Lu, Jiahao Wang, Ye Feng, Ping Luo, and Ying Shan. Llama pro: Progressive llama with block expansion. arXiv preprint arXiv:2401.02415, 2024.
  • [33] Noam Shazeer. Glu variants improve transformer. arXiv preprint arXiv:2002.05202, 2020.
  • [34] Zoltan Csaki, Pian Pawakapan, Urmish Thakker, and Qiantong Xu. Efficiently adapting pretrained language models to new languages. arXiv preprint arXiv:2311.05741, 2023.
  • [35] Aleksandr Nikolich, Konstantin Korolev, and Artem Shelmanov. Vikhr: The family of open-source instruction-tuned large language models for russian. arXiv preprint arXiv:2405.13929, 2024.
  • [36] Ali Farghaly and Khaled Shaalan. Arabic natural language processing: Challenges and solutions. ACM Transactions on Asian Language Information Processing (TALIP), 8(4):1–22, 2009.
  • [37] Sabri Elkateb, William J Black, Piek Vossen, David Farwell, Horacio Rodriguez, Adam Pease, Musa Alkhalifa, and Christiane Fellbaum. Arabic wordnet and the challenges of arabic. In Proceedings of the International Conference on the Challenge of Arabic for NLP/MT, pages 15–24, 2006.
  • [38] Iyad Issa. Morphological complexity in arabic spelling and its implication for cognitive processing. Journal of Psycholinguistic Research, 52(1):331–357, 2023.
  • [39] Taku Kudo and John Richardson. Sentencepiece: A simple and language independent subword tokenizer and detokenizer for neural text processing. arxiv 2018. arXiv preprint arXiv:1808.06226, 1808.
  • [40] Andrey Gromov, Kushal Tirumala, Hassan Shapourian, Paolo Glorioso, and Daniel A Roberts. The unreasonable ineffectiveness of the deeper layers. arXiv preprint arXiv:2403.17887, 2024.
  • [41] Wissam Antoun, Fady Baly, and Hazem Hajj. Arabert: Transformer-based model for arabic language understanding. arXiv preprint arXiv:2003.00104, 2020.
  • [42] Chaofan Tao, Qian Liu, Longxu Dou, Niklas Muennighoff, Zhongwei Wan, Ping Luo, Min Lin, and Ngai Wong. Scaling laws with vocabulary: Larger models deserve larger vocabularies. arXiv preprint arXiv:2407.13623, 2024.
  • [43] Kaj Bostrom and Greg Durrett. Byte pair encoding is suboptimal for language model pretraining. arXiv preprint arXiv:2004.03720, 2020.
  • [44] Ali Elfilali, Hamza Alobeidli, Clémentine Fourrier, Basma El Amel Boussaha, Ruxandra Cojocaru, Nathan Habib, and Hakim Hacid. Open arabic llm leaderboard. https://huggingface.co/spaces/OALL/Open-Arabic-LLM-Leaderboard, 2024.

### v2

@@ Abstract
11footnotetext: Equal contributions22footnotetext: Formerly affiliated with misraj.ai at the time of this research; m.omar.hadid.1997@gmail.com 33footnotetext: Kuwain: \<كُوَيْن>: is a diminutive form of the Arabic word (Kawn), which means "universe". So, "Kuwain" means "tiny universe" or "little cosmos".Kuwain is one of a series of Arabic-English multilingual Large Language Models (LLMs), embodying the concept of a vast amount of knowledge condensed into a compact, accessible form, much like a miniature universe of information.
@@ 1 Introduction
Figure 1: Model performance vs size visualization: Points represent models on the Arabic NLP benchmark, where point size indicates efficiency (average score divided by model size). Larger points show better performance-to-size ratio.
@@ 4 Method
Figure 2: (a) Vocabulary expansion by adding new Arabic tokens to the tokenizer. (b) Extension of model layers, where newly added layers were trainable while the original layers remained frozen.
@@ 4.1 Extending Model’s Layers
Figure 3: models decoder architecture.
@@ 5.1 Number and Placement of New Layers
Figure 4: Training loss charts for Kuwain under different configurations: (a) Freezing the last layer leads to instability, (b) Enabling the last layer and avoiding consecutive layers yields stable training, (c) Stacking new layers consecutively results in degraded performance.
@@ 5.2 Tokenizer Comparison
Table 1: Comparison of tokenization metrics across different Arabic language models. The vocab size used presents the Arabic vocab for each model, where the full vocab has a different number.
@@ 5.3 Data Proportion
Table 2: Performance comparison on English benchmarks between TinyLlama and two variants of Kuwain: trained with 20% English data (Kuwain) and with less than 20% English data (Kuwain-lt-$\phi$). Results show how English performance varies with data proportion.
@@ 5.4 Conventional Training (Without Layer Extension)
Table 3: English evaluation of TinyLlama vs. Kuwain-Naive (standard pre-training).
@@
Table 4: Arabic evaluation of Kuwain-Naive vs. our model Kuwain.
@@ 6 Evaluation
Table 5: Arabic Benchmark Evaluation of Kuwain and TinyLlama
@@
Table 6: Arabic Leader board evaluation